OpenStudy (dillonmerv):

Let u =⟨2,−5⟩, v =⟨−1,1⟩, and w =⟨−2,4⟩. Find the vector x that satisfies 2u −v +x =10x +w.

OpenStudy (sshayer):

2u=(4,10) Let x=(a,b) 2u-v+x=(4,-10)-(-1,1)+(a,b)=(4+1+a,-10-1+b) 10x+w=(10a,10b)+(-2,4)=(10a-2,10b+4) now equate and find a,b

OpenStudy (loser66):

is it not that 2u-v +x= 10x+w, then 9x= 2u-v-w? and then x= (1/9) (2u-v-w) ?